# KeyMPs: One-Shot Vision-Language Guided Motion Generation by Sequencing DMPs for Occlusion-Rich Tasks

### v1

@@ VI REAL ROBOT EXPERIMENT
Building on the experimental setup detailed in §V, we further validated the effectiveness of the KeyMPs framework in a real-world setting. This experiment focused on assessing whether the generated executable motions faithfully align with the intent expressed in the multimodal input, as well as on the framework’s performance under practical conditions.
@@ VI-A Research Questions
Figure 15: Object cutting environment in real-world setting.
@@ VI-B1 Task Description
We designed ten task cases (see Table IV) to assess the performance of our framework in a real-world setting. Each task was executed, and the outcomes were qualitatively evaluated on the basis of successful completion of the cutting tasks as per the user’s input.
TABLE IV: List of prepared cutting tasks in the real environment
@@
Figure 16: Results of real robot experiment for all task cases in Table IV. Red lines in the first column represent the visualization of the VLMs generated keypoint pairs with green numbers for the order of cuts. In the last column, the yellow-dotted lines represent how the objects are cut.
@@ VI-B2 Primitive Dictionary Preparation
Instead of employing a sawing motion, the forward cutting primitive was utilized to minimize object displacement when performing the cutting motion. Visualizations of primitives used in the real experiments are presented in Appendix D.
VI-C Experiment 4: Feasibility in a Real-World Setting
@@ VII DISCUSSION
The results of the evaluation confirm that our KeyMPs framework generates complex executable motions in on e shot that closely align with the intent expressed in the multimodal input, while also achieving consistent and data-efficient DMP-based motion generation in an occlusion-rich task. Moreover, an additional experiment detailed in Appendix E demonstrates that integrating multimodal input significantly outperforms using language input alone, even when supplemented with additional context. By effectively handling learned and spatial scaling parameters separately and leveraging VLMs’ conceptual understanding and high-level reasoning capabilities, the framework requires only a single demonstration for each primitive.
@@ VIII CONCLUSION
This paper introduced KeyMPs, a novel framework that enhances robotic motion generation through the utilization of VLMs and sequencing DMPs. By leveraging VLMs’ high-level reasoning for selecting reference DMPs and their spatial awareness for keypoint pairs generation, KeyMPs effectively bridges different high-level language instructions with motion generation through sequencing DMPs. The structured decomposition for sequencing multiple DMPs enables one-shot vision-language guided motion generation that is adaptable and generalizable in occlusion-rich tasks while reducing the need for extensive human demonstrations. Validated through both simulation and real-world experiments, KeyMPs consistently produced motions that accurately aligned with the intent expressed in the multimodal input, achieving a high degree of consistency and data efficiency.
@@
VIII-A Pixel-Based Object Detection
@@
VIII-B Transforming 2D Keypoint Pairs from Local to Global Coordinates
@@
    Normalization: Normalize the translated keypoint $\mathbf{p}^{\prime}$ to a $[0,1]$ range based on the image size:
@@
VIII-C Integrating Height Information into Keypoint Pairs
@@
    Both Keypoints Integrate Object’s Height
@@
    Only Starting Keypoint Integrates Object’s Height
@@
    Only Ending Keypoint Integrates Object’s Height
@@
VIII-D Visualizations of the Primitives Used in the Experiments
This appendix offers visual representations of the primitives used in our experiments, which designed to facilitate specific actions within the simulation and real-world robot tasks. Visualizations of the primitives used in the simulation experiments are presented in Fig. 17, while visualizations of the primitives used in the real robot environment are presented in Fig. 18.
@@ VIII-D Visualizations of the Primitives Used in the Experiments
VIII-E Ablation: Necessity of Image Input
VIII-E1 Objective
@@
VIII-E2 Comparison Methods
TABLE V: Information contained in input for each method
@@ VIII-E2 Comparison Methods
We evaluated our framework against ablation methods that employed LLMs instead of VLMs, referred to as KeyMPs-text. The prompts for KeyMPs-text depended only on linguistic input containing both the name of the object together with its shape. Table V summarizes the information received by each method.
VIII-E3 Results
@@
TABLE VI: Success rates of keypoint pair designs in experiment 3

### v2

@@ V-F3 Results
V-G Experiment 4: Evaluating Framework Flexibility for Broader Applications
V-G1 Objective
To answer (P4), we test the flexibility of our framework by applying it to a new application: cake icing. This task was chosen to assess whether the core principles of KeyMPs’ VLM driven keypoint generation combined with a single motion primitive can generalize to a broader variety of creative and intricate tasks, with scenarios listed in Table 4.
V-G2 Results
TABLE IV: List of prepared cake icing tasks in the real environment
Case Object Input Prompt
1 Round cake Put 15 icings around the cake, provide some space from the cake edge and between each icing.
2 Round cake Make 2 layers of icing around the cake, provide some space from the edges and between layers.
3 Round cake Combine 2 types of designs, icing dots around the cake and a spiral icing line design.
4 Round cake I’m having my 27th birthday, put the digits on the cake using seven-segment characters as icing lines.
5 Square cake Draw a 5-pointed star using lines in the middle of the cake.
6 Round cake Create a simple interesting design.
7 Square cake Create a simple interesting design.
8 Square cake I want to put 1 dot of icing on the middle of each of the 10 slices.
As shown in Fig. 15, the framework demonstrated significant flexibility across all scenarios. The system generated simple geometric patterns (Cases 1, 2), combined different design elements (Case 3), and produced motions from abstract concepts like numbers and shapes (Cases 4, 5).
Furthermore, the framework handled ambiguous prompts by generating creative yet context-aware designs (Cases 6, 7). For example, when tasked with creating a ”simple interesting design,” it produced a circular pattern for the round cake and a square-based pattern for the square cake. The framework also solved multi-step reasoning problems, such as dividing a cake before placing icing (Case 8).
Collectively, these outcomes show that the KeyMPs architecture has the potential for broader application. By simply providing a new primitive and leveraging the VLMs’ contextual understanding, the framework successfully generates motions for this distinct application. This directly answers (P4) by demonstrating the framework’s capacity for generalization to new tasks.
Figure 15: Results of the cake icing experiments, demonstrating the generalizability of the KeyMPs framework. Right-most column shows the final outcome for the corresponding task case in Table IV.
@@ VI REAL ROBOT EXPERIMENT
Building on the experimental setup detailed in §V-B1, we further validated the effectiveness of the KeyMPs framework in a real-world setting. This experiment assessed if the generated motions align with the multimodal input intent and evaluated the framework’s practical performance.
@@ VI-B1 Task Description
We designed ten task cases (see Table V) to assess the performance of our framework in a real-world setting. Each task was executed, and the outcomes were qualitatively evaluated on the basis of successful completion of the cutting tasks as per the user’s input.
Figure 16: Object cutting environment in real-world setting.
TABLE V: List of prepared cutting tasks in the real environment
@@ VI-B2 Primitive Dictionary Preparation
Initially, a sawing primitive used in §V was utilized. However, due to displacement resulting from this motion, the primitive was simplified to a forward cutting primitive to mitigate displacement. Visualizations of primitives used in the real experiments are presented in Appendix D.
VI-C Experiment 5: Feasibility in a Real-World Setting
@@ VI-C1 Objective
Figure 17: Results of real robot experiment for all task cases in Table V. Red lines in the first column represent the visualization of the VLMs generated keypoint pairs with green numbers for the order of cuts. In the last column, the yellow-dotted lines represent how the objects are cut.
@@ VII DISCUSSION
The results of the evaluation confirm that our KeyMPs framework generates complex executable motions in one shot that closely align with the intent expressed in the multimodal input, while also achieving consistent and data-efficient DMP-based motion generation in occlusion-rich tasks. Moreover, an additional experiment detailed in Appendix F demonstrates that integrating multimodal input significantly outperforms using language input alone, even when supplemented with additional context. By effectively handling learned and spatial scaling parameters separately and leveraging VLMs’ conceptual understanding and high-level reasoning capabilities, the framework requires only a single demonstration for each primitive.
@@ VIII CONCLUSION
This paper introduced KeyMPs, a novel framework that enhances robotic motion generation through the utilization of VLMs and sequencing DMPs. By leveraging VLMs’ high-level reasoning for selecting reference DMPs and their spatial awareness for keypoint pairs generation, KeyMPs effectively bridges different high-level language instructions with motion generation through sequencing DMPs. The structured decomposition for sequencing multiple DMPs enables one-shot vision-language guided motion generation that is adaptable and generalizable in occlusion-rich tasks while reducing the need for extensive human demonstrations. Validated through both simulation and real-world experiments, KeyMPs consistently produced motions that accurately aligned with the intent expressed in the multimodal input, achieving a high degree of consistency and data efficiency in multiple tasks.
@@
A. Pixel-Based Object Detection
@@
B. Transforming 2D Keypoint Pairs from Local to Global Coordinates
@@
    Normalization: Normalize the translated keypoint $\mathbf{p}^{\prime}$ to a [0,1][0,1][ 0 , 1 ] range based on the image size:
@@
C. Integrating Height Information into Keypoint Pairs
@@
    both keypoints integrate object’s height,
@@
    only starting keypoint integrates object’s height, and
@@
    only ending keypoint integrates object’s height.
@@
Figure 20: 3D projection of the icing line primitive used in simulation. Time progression is represented by the change in color from red to green to blue.
Figure 21: Visualization of the user survey results conducted to validate our predefined ground-truth lines. The heatmap shows the aggregated drawings from five participants for various cutting tasks. The intensity of the red color indicates the frequency of drawn lines in that area, with stronger red signifying higher agreement among subjects. The analysis confirmed that our established ground-truth lines align well with these high-frequency areas.
D. Visualizations of the Primitives Used in the Experiments
This appendix offers visual representations of the primitives used in our experiments, designed to facilitate specific actions within the simulation and real-world robot tasks. Visualizations of the primitives for the simulation experiments are presented in Fig. 18 for the cutting task and Fig. 20 for the icing task, while the primitive used in the real robot environment is presented in Fig. 19.
E. Object Cutting Task Ground Truth Lines Survey Results
To confirm that the ground-truth lines we established from domain knowledge were consistent with general user expectations, we conducted a validation survey with five consenting subjects (ages: 24-28). The survey procedure for each object cutting task case, as detailed in Table I, consisted of providing participants with an image of the object, an indicator of its real-world size, and the corresponding text prompt. The subjects were then instructed to draw the cutting lines they would expect a robot to perform directly onto the image using a marker.
The collected drawings were aggregated and visualized as a heatmap in Fig. 21. In this visualization, the intensity of the red color corresponds to the frequency with which participants drew a line in a specific area, indicating spatial agreement. This analysis confirmed that our pre-established ground-truth lines are consistent with the most commonly agreed-upon cutting patterns identified in the survey, validating our initial expert-based approach.
F. Ablation: Necessity of Image Input
VIII-A Objective
@@
VIII-B Comparison Methods
TABLE VI: Comparison of input modalities for each method in the ablation study on the importance of image input
@@ VIII-B Comparison Methods
We evaluated our framework against ablation methods that employed LLMs instead of VLMs, referred to as KeyMPs-text. The prompts for KeyMPs-text depended only on linguistic input containing both the name of the object together with its shape. Table VI summarizes the information received by each method.
VIII-C Results
@@
[…]
TABLE VII: Success rates for keypoint pair generation in the image input ablation study